The length of a rectangle is 6ft longer than its width. If the perimeter of the rectangle is 40ft , find its area.
Answer by london maths tutor(243)

Let's the length be L and width be w
L = 6 + w
Perimeter of the rectangle = 2(L+w)=40
2(6 + w + w) = 40
2(6 + 2w) = 40
12 + 4w = 40
4w = 40 - 12
4w = 28
w = 28/4
w = 7 ft
Therefore, L = 6 + 7 = 13 ft
Area = L * w = 13*7 = 91ft^2